Solve the linear equation given the following:
(1/2, 1) ; (-1,-1/2)

Answer by stanbon(75887)

Solve the linear equation given the following:
(1/2, 1) ; (-1,-1/2)
------
slope = (1--(1/2)/(1/2 - -1) = (3/2)/(3/2) = 1
------
Form: y = mx + b
Solve for "b":
1 = 1*(1/2) + b
b = 1/2
------
Equation:
y = x + 1/2
=============
Cheers,
Stan H.
==========